# On orthogonality graphs of Okubo algebras

### v1

@@ 5 The case of a field containing a primitive cube root of unity
In this section we pay special attention to the orthogonality graph of an Okubo algebra with nonzero idempotents in the case of a field $\mathbb{F}$, $\operatorname{char}\mathbb{F}\neq 3$, which contains a primitive cube root of unity $\omega$. Recall that, by Proposition 2.12(2), such an Okubo algebra is always isomorphic to the pseudo-octonion algebra $P_{8}(\mathbb{F})$, that is, it can be considered as the set ${\mathfrak{sl}}_{3}(\mathbb{F})$ of traceless $3\times 3$ matrices over $\mathbb{F}$ with the multiplication given by Eq. (2.1). In particular, if the field $\mathbb{F}$ is algebraically closed, then by definition any Okubo algebra over $\mathbb{F}$ is isomorphic to $P_{8}(\mathbb{F})$, so it always contains a nonzero idempotent (see also [9, Theorem 2.8] which shows that any finite-dimensional flexible composition algebra over an algebraically closed field with $\operatorname{char}\mathbb{F}\neq 2,3$ contains a nonzero idempotent).

### v2

@@ 5 The case of a field containing a primitive cube root of unity
In this section we pay special attention to the orthogonality graph of an Okubo algebra with nonzero idempotents in the case of a field $\mathbb{F}$, $\operatorname{char}\mathbb{F}\neq 3$, which contains a primitive cube root of unity $\omega$. Recall that, by Proposition 2.12(2), such an Okubo algebra is always isomorphic to the pseudo-octonion algebra $P_{8}(\mathbb{F})$, that is, it can be considered as the set ${\mathfrak{sl}}_{3}(\mathbb{F})$ of traceless $3\times 3$ matrices over $\mathbb{F}$ with the multiplication given by Eq. (2.1). In particular, if the field $\mathbb{F}$ is algebraically closed, then by definition any Okubo algebra over $\mathbb{F}$ is isomorphic to $P_{8}(\mathbb{F})$, so it always contains a nonzero idempotent (see also [6, Theorem 2.8] which shows that any finite-dimensional flexible composition algebra over an algebraically closed field with $\operatorname{char}\mathbb{F}\neq 2,3$ contains a nonzero idempotent).